OpenStudy (anonymous):

Simplify by taking roots of the numerator and the denominator. ^3sqrt125/27

OpenStudy (anonymous):

We can solve by distributing the cube root to the numerator and denominator separately. \[\sqrt[3]{125/27} = \sqrt[3]{125}/\sqrt[3]{27}\]

OpenStudy (anonymous):

okay. I'm confused on how to do this. :)

OpenStudy (anonymous):

The question to answer is now x*x*x = 125. Find x. Then z*z*z = 27. Find z. Your answer will then your simplified answer will be x/z.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

5*5*5=125

OpenStudy (anonymous):

3*3*3

OpenStudy (anonymous):

Correct. So what is the final simplified answer?

OpenStudy (anonymous):

5^3/3^3?

OpenStudy (anonymous):

Not quite. We don't need it raised to the third power because we took the cube root. \[\sqrt[3]{125/27} = 5/3\]

OpenStudy (anonymous):

oooh ok! :) Thank you :)